# Towards High-Quality and Efficient Speech Bandwidth Extension with Parallel Amplitude and Phase Prediction

### v1

@@ V-C Analysis and Discussion
TABLE VI: Experimental Results for the Ablation Studies with Input Sampling Rate of 8 kHz and Target Sampling Rate of 48 kHz
@@
V-C1 Ablation Studies
We implemented ablation studies on discriminators and between-stream connections to investigate the roles of each discriminator and the effects of interactions between amplitude stream and phase stream. All the experiments were conducted with the input sampling rate of 8 kHz and target sampling rate of 48 kHz, and the experimental results are depicted in Table VI. Due to the minimal phase differences in the high-frequency components for the BWE targeting 48 kHz, we calculated the AWPD metric only in the frequency band of 4 kHz to 8 kHz while calculating the LSD metric on the whole frequency band. We further visualized the spectrograms of the natural wideband 48 kHz speech waveform and speech waveforms generated by the ablation models of AP-BWE on discriminator, as illustrated in Fig. 5.
As shown in the upper half of Table VI, for the ablation of discriminators, all the discriminators contributed to the overall performance of our proposed AP-BWE. We first ablated the MPD to solely train the AP-BWE model only with discriminators at the spectral level. Although there is only a slight decrease in all metrics, the spectrogram (AP-BWE w/o MPD) reveals a smearing effect along the frequency axis as shown in Fig. 5, resulting in a perceptible harshness in the extended speech. Subsequently, in our preliminary experiments, we separately ablated MRPD and MRAD. In both cases, the metrics showed only slight decreases, and the spectrograms appeared normal. However, when we simultaneously ablated both of them (AP-BWE w/o MRDs), although the metrics still decreased insignificantly, a noticeable over-smoothing could be observed in the frequency band from 12 kHz to 24 kHz of the spectrogram. This is because with the sole utilization of MPD, the minimum period of its sub-discriminator was 2, so the frequency band range it can discriminate was only from 0 to 12 kHz. As this over-smoothing phenomenon is present in the high-frequency range, it has not had a substantial impact on the perceived quality of extended speech. We further ablated all discriminators (AP-BWE w/o Disc.), the LSD metric experienced a significant decline, and the extended portions across the entire spectrogram exhibited severe over-smoothing, greatly compromising the quality of the extended speech. This indicated that the training strategy of GAN was indispensable for the current AP-BWE model.
Moreover, we ablated the between-stream connections. As shown in the last row of Table VI, the information interaction between the amplitude stream and the phase stream did contribute to the quality of the extended speech. To investigate the influence of one stream on another, we selectively ablated each of the connections. Our observations revealed that when ablating the connection from the amplitude stream to the phase stream (A $\rightarrow$ P), the AWPD metrics exhibited a deterioration compared to ablating the connection from the phase stream to the amplitude stream (P $\rightarrow$ A), and there was also a decrease in ViSQOL, indicating that amplitude information played a role in the modeling of phase. This conclusion aligns with our previous work [34], where the phase spectrum can be predicted from the amplitude spectrum.
TABLE VII: Experimental Results for the Cross-Dataset Evaluation on the Libri-TTS and HiFi-TTS datasets
Libri-TTS (8 kHz normal-$\rightarrow$ 24 kHz)
@@ V-C2 Cross-Dataset Validation
Since the speech data in a corpus is recorded in a fixed environment, models trained exclusively on a single corpus may adapt to the specific characteristics of the recording environment. To evaluate the models’ generalization abilities across different corpora, we conducted cross-dataset experiments on models trained with the input and target sampling rates of 8 kHz and 48 kHz, respectively. We selected two high-quality datasets, namely Libri-TTS [62] and HiFi-TTS [63]. The Libri-TTS dataset consists of 585 hours of speech data at a 24 kHz sampling rate. For evaluation, we exclusively utilized the “test-clean” set, containing 4,837 audio clips. The HiFi-TTS dataset contains about 292 hours of speech from 10 speakers with at least 17 hours per speaker sampled at 44.1 kHz, and we also only evaluated the models on its test set which contains 1000 audio clips. The experimental results are depicted in Table VII, where the LSD scores were computed by downsampling all the extended speech waveforms from 48 kHz to the original sampling rates of the datasets, the AWPD metrics were calculated only in the 4kHz $\sim$ 8 kHz frequency band for a more intuitive comparison, and the ViSQOL scores were computed by upsampling all the speech waveforms to 48 kHz.
For the evaluation on the Libri-TTS dataset, as depicted in the upper half in Table VII, our proposed AP-BWE still achieved the best performance among all the metrics. For NU-Wave2 and UDM+, their performance on the Libri-TTS dataset was noticeably degraded compared to VCTK. This indicated a strong dependency of waveform-based methods on the training corpus, whereas spectrum-based approaches, by capturing temporal and spectral characteristics from the waveforms, exhibited adaptability to various data recording environments. The evaluation results on the HiFi-TTS dataset are depicted in the lower half in Table VII. Spectrum-based methods consistently outperform waveform-based methods, particularly in terms of the ViSQOL metric. Unsurprisingly, mdctGAN achieved the best performance since it was trained on the combination of the VCTK and HiFi-TTS datasets. However, our AP-BWE, trained solely on VCTK, still achieved a comparable performance and even surpassed mdctGAN in the AWPD metrics. This suggests that our phase prediction exhibited sufficient generalization across different corpora. The overall performance of the models on the HiFi-TTS dataset was not as impressive as on Libri-TTS. This discrepancy could be attributed to the fact that, despite the HiFi-TTS dataset having a sampling rate of 44.1 kHz, its effective bandwidth falls between 13 kHz and 22.05 kHz. In contrast, the waveforms generated by the BWE models cover the entire bandwidth. As a result, there may be mismatches in the calculation of metrics.
@@ VI Conclusion
In this paper, we introduced AP-BWE, a GAN-based BWE model that can efficiently achieve high-quality wideband waveform generation. The generator of AP-BWE performed direct recovery of high-frequency amplitude and phase information from the narrowband amplitude and phase spectra through an all-convolutional structure and all-frame-level operations, significantly enhancing generation efficiency. Moreover, multiple discriminators applied on the time-domain waveform, amplitude spectrum, and phase spectrum noticeably elevated the overall generation quality. The major contribution of the AP-BWE lay in the direct extension of the phase spectrum. This allowed for the precise modeling and optimization of both the amplitude and phase spectra simultaneously, significantly enhancing the quality of the extended speech without compromising the trade-offs between the two. Experimental results on the VCTK-0.92 dataset showcased that our proposed AP-BWE achieved SOTA performance for tasks with target sampling rates of both 16 kHz and 48 kHz. Spectrogram visualizations underscored the robust capability of our model in recovering high-frequency harmonic structures, effectively enhancing the intelligibility of speech signals, even in scenarios with extremely low input speech bandwidth. In future work, our AP-BWE model can be further applied to assist generative models trained on low-sampling-rate datasets to improve their synthesized speech quality.

### v2

@@ V-C Analysis and Discussion
V-C1 Ablation Studies
We implemented ablation studies on discriminators and between-stream connections to investigate the roles of each discriminator and the effects of interactions between amplitude stream and phase stream. All the experiments were conducted with the source sampling rate of 8 kHz and target sampling rate of 48 kHz, and the experimental results are depicted in Table VIII. Due to the minimal phase differences in the high-frequency components for the BWE targeting 48 kHz, we calculated the AWPD metric only in the frequency band of 4 kHz to 8 kHz while calculating the LSD metric on the whole frequency band. We further visualized the spectrograms of the natural wideband 48 kHz speech waveform and speech waveforms generated by the ablation models of AP-BWE on discriminator, as illustrated in Fig. 6.
TABLE VIII: Experimental Results for the Ablation Studies with Source Sampling Rate of 8 kHz and Target Sampling Rate of 48 kHz
@@ V-C1 Ablation Studies
As shown in the upper half of Table VIII, for the ablation of discriminators, all the discriminators contributed to the overall performance of our proposed AP-BWE. We first ablated the MPD to train the AP-BWE model solely with discriminators at the spectral level. Although there is only a slight decrease in all metrics, the spectrogram (AP-BWE w/o MPD) reveals a smearing effect along the frequency axis as shown in Fig. 6, resulting in a perceptible harshness in the extended speech. Subsequently, in our preliminary experiments, we separately ablated MRPD and MRAD. In both cases, the metrics showed only slight decreases, and the spectrograms appeared normal. However, when we simultaneously ablated both of them (AP-BWE w/o MRDs), although the metrics still decreased insignificantly, a noticeable over-smoothing could be observed in the frequency band from 12 kHz to 24 kHz of the spectrogram. This is because with the sole utilization of MPD, the minimum period of its sub-discriminator was 2, so the frequency band range it can discriminate was only from 0 to 12 kHz. As this over-smoothing phenomenon is present in the high-frequency range, it has not had a substantial impact on the perceived quality of extended speech. We further ablated all discriminators (AP-BWE w/o Disc.), the LSD metric experienced a significant decline, and the extended portions across the entire spectrogram exhibited severe over-smoothing, greatly compromising the quality of the extended speech. This indicated that the training strategy of GAN was indispensable for the current AP-BWE model.
Moreover, we ablated the between-stream connections. As shown in the last row of Table VIII, the information interaction between the amplitude stream and the phase stream did contribute to the quality of the extended speech. To investigate the influence of one stream on another, we selectively ablated each of the connections. Our observations revealed that when ablating the connection from the amplitude stream to the phase stream (A $\rightarrow$ P), the AWPD metrics exhibited a deterioration compared to ablating the connection from the phase stream to the amplitude stream (P $\rightarrow$ A), and there was also a decrease in ViSQOL, indicating that amplitude information played a role in the modeling of phase. This conclusion aligns with our previous work [34], where the phase spectrum can be predicted from the amplitude spectrum.
TABLE IX: Experimental Results for the Cross-Dataset Evaluation on the Libri-TTS and HiFi-TTS Datasets
Libri-TTS (8 kHz $\rightarrow$ 24 kHz)
@@ V-C2 Cross-Dataset Validation
Since the speech data in a corpus is recorded in a fixed environment, models trained exclusively on a single corpus may adapt to the specific characteristics of the recording environment. To evaluate the models’ generalization abilities across different corpora, we conducted cross-dataset experiments on models trained with the source and target sampling rates of 8 kHz and 48 kHz, respectively. We selected two high-quality datasets, namely Libri-TTS [63] and HiFi-TTS [64]. The Libri-TTS dataset consists of 585 hours of speech data at a 24 kHz sampling rate. For evaluation, we exclusively utilized the “test-clean” set, containing 4,837 audio clips. The HiFi-TTS dataset contains about 292 hours of speech from 10 speakers with at least 17 hours per speaker sampled at 44.1 kHz, and we also only evaluated the models on its test set which contains 1000 audio clips. The experimental results are depicted in Table IX, where the LSD scores were computed by downsampling all the extended speech waveforms from 48 kHz to the original sampling rates of the datasets, the AWPD metrics were calculated only in the 4kHz $\sim$ 8 kHz frequency band for a more intuitive comparison, and the ViSQOL scores were computed by upsampling all the speech waveforms to 48 kHz.
For the evaluation on the Libri-TTS dataset, as depicted in the upper half in Table IX, our proposed AP-BWE still achieved the best performance among all the metrics. For NU-Wave2 and UDM+, their performance on the Libri-TTS dataset was noticeably degraded compared to VCTK. This indicated a strong dependency of waveform-based methods on the training corpus, whereas spectrum-based approaches, by capturing temporal and spectral characteristics from the waveforms, exhibited adaptability to various data recording environments. The evaluation results on the HiFi-TTS dataset are depicted in the lower half in Table IX. Compared to the waveform-based methods, spectrum-based methods still outperform in terms of the overall speech quality as indicated by the ViSQOL metric. Compared to NU-Wave2 and UDM+, the advantage of mdctGAN on the HiFi-TTS test set is far less pronounced than on the Libri-TTS test set, especially in terms of the LSD metric. This suggests that different models exhibit varying generalization abilities across different datasets. However, our proposed AP-BWE still shows a significant advantage over other baseline models across all metrics, further demonstrating the superior generalization ability of our model.
@@ VI Conclusion
In this paper, we introduced AP-BWE, a GAN-based BWE model that can efficiently achieve high-quality wideband waveform generation. The generator of AP-BWE performed direct recovery of high-frequency amplitude and phase information from the narrowband amplitude and phase spectra through an all-convolutional structure and all-frame-level operations, significantly enhancing generation efficiency. Moreover, multiple discriminators applied on the time-domain waveform, amplitude spectrum, and phase spectrum noticeably elevated the overall generation quality. The major contribution of the AP-BWE lay in the direct extension of the phase spectrum. This allowed for the precise modeling and optimization of both the amplitude and phase spectra simultaneously, significantly enhancing the quality of the extended speech without compromising the trade-offs between the two. Experimental results on the VCTK-0.92 dataset showcased that our proposed AP-BWE achieved SOTA performance for tasks with target sampling rates of both 16 kHz and 48 kHz. Spectrogram visualizations underscored the robust capability of our model in recovering high-frequency harmonic structures, effectively enhancing the intelligibility of speech signals, even in scenarios with extremely low source speech bandwidth. In future work, our AP-BWE model can be further applied to assist generative models trained on low-sampling-rate datasets in improving their synthesized speech quality.